Suppose the required reserve ratio is 0.20. Total bank deposits are $200 million and the bank holds $50 million in reserves. How much more money could the bank create if it does not hold excess reserves?

A. $5 million B. $25 million C. $30 million D. $50 million
Answer #1

Answer: D. $50 million

Required reserve = Deposits * Required reserve ratio

Required reserve = $200 * 0.20 = $40

Given: Bank holds $50 million in reserves.

Excess reserve = Total reserve - Required reserve

Excess reserve = $50 - $40

Excess reserve = $10 million

Money creation = Money multiplier * Excess reserve

Money multiplier = 1/Required reserve ratio

Money creation = (1/0.20) * 10 = 5 * 10 = 50

Banks would create $50 million, if it does not hold excess reserves.